# SalUn: Empowering Machine Unlearning via Gradient-based Weight Saliency in Both Image Classification and Generation

### 06 Nov 2023 v2

@@ 7 Conclusion
Recognizing the shortcomings of existing MU (machine unlearning) approaches, we introduce the innovative concept of weight saliency in MU, leading to the SalUn framework. The proposed saliency unlearning has showcased its effectiveness in addressing the limitations of current MU methods, and applying to both image classification and generation tasks. As a notable application, we show that SalUn stays effective in preventing stable diffusion from generating images with harmful content, even when given inappropriate image prompts. As other future research directions, we will explore the theoretical foundations of weight saliency-aided unlearning and consider other MU applications in generative AI, such as large language models.
@@ 8 Acknowledgement
This work was supported by the Cisco Research Award and partially supported by the NSF Grant IIS-2207052.
@@ Appendix A Soft-thresholding SalUn
where $i=1$ (or $2$) corresponds to the objective function of SalUn for classification (6) (or generation (7)), and $\beta>0$ is a regularization parameter. Note that unlike (6) and (7), the objective function of SalUn is defined over the entire model weights $\bm{\theta}$} since there is no weight saliency map known in advance, i.e., letting $\mathbf{m}_{\mathrm{S}}=\mathbf{1}$ in (4). Problem (8) can be efficiently solved using the proximal gradient algorithm (Parikh et al., 2014), wherein the $\ell_{1}$ penalty term is handled efficiently through a closed-form proximal operation known as soft-thresholding. In contrast to the hard-thresholding implementation of SalUn, the soft-thresholding approach requires tuning an additional hyperparameter $\beta$, with the trade-off of not resorting to the hard threshold $\gamma$ in (3). In practice, we find that employing a linearly decaying scheduler for $\beta$ (i.e., promoting weight saliency sparsity during the early unlearning stages) yields better results than other schemes, approaching the performance of SalUn based on hard thresholding.
@@ B.1 Additional training and unlearning settings
For the Retrain method, training is conducted over $182$ epochs using the SGD optimizer with a cosine-scheduled learning rate initialized at $0.1$. For both FT and RL, training spans 10 epochs with a fixed learning rate of $0.1$. GA’s training settings involve a 5-epoch learning rate search within the interval $10^{-6},10^{-4}]$. In the case of IU, we explore the parameter $\alpha$ associated with the woodfisher Hessian Inverse approximation within the range $[1,20]$. For $\ell_{1}$-sparse, a learning rate search for the parameter $\gamma$ is executed within the range $10^{-6},10^{-4}]$, while maintaining a constant learning rate of 0.1. For the BS method, the FGSM step size is defined as 0.1. Both BS and BE methods involve a 10-epoch learning rate search in the interval $10^{-6},10^{-4}]$. Lastly, for SalUn and SalUn-soft, training is carried out over 10 epochs with a set learning rate of 0.1.
@@
Refer to caption Refer to caption Refer to caption (a) UA for classification (b) MIA for classification (c) TA for classification
Figure A1: Performance of SalUn with different saliency sparsity vs Retrain for classification. The settings follow Table 1. Points above Retrain indicate over-forgetting, while points below Retrain represent under-forgetting.
Ablation study on sparsity choice of weight saliency.
Recall from (3) that the choice of the weight saliency sparsity threshold could be a crucial hyperparameter in our approach. In our experiments, we have set the default sparsity threshold to 50%. Fig. A1 and Fig. A2 provide a more detailed examination of the sparsity ratio. Specifically, Fig. A1-(a), (b) and (c) present the performance of random data forgetting in image classification using ResNet-18 on CIFAR-10 (with the experiment setup same as Table 1). Fig. A2-(a) and (b) present the performance of class-wise forgetting for image generation using DDPM on CIFAR-10 with the same setting as Fig 4. In both cases, the performance of Retrain is provided for comparison.
Refer to caption Refer to caption (a) UA vs. sparsity (b) FID vs. saprsity
Figure A2: Performance of SalUn with different saliency sparsity vs Retrain in the tas of image generation. The settings follow Fig 4.
In the context of MU for image classification, choosing a 50% sparsity threshold is a reasonable option. A higher sparsity of the weight saliency may result in under-forgetting, as evidenced by the lower unlearning accuracy (Jia et al., 2023) (namely, the higher classification accuracy on the forgetting data points) and the lower MIA (Jia et al., 2023) (namely, making it challenging to infer the forgetting identity of a training point), compared to Retrain. Moreover, a higher saliency sparsity yields better testing accuracy. This is not surprising since as shown in (4), the higher saliency sparsity indicates fewer weights of the pre-trained model to be modified during unlearning. Conversely, selecting a lower sparsity ratio may result in over-forgetting due to the excessive number of weights to be modified. In the context of MU for image generation (Fig. A2), a higher sparsity ratio also leads to under-forgetting, while a much lower sparsity ratio causes a rise in FID, introduced by over-forgetting.
@@
Figure A3: Image generations of SalUn with different saliency sparsity. The settings follow Fig 4.
Figure A4: The relationship between optimal saliency sparsity and forgetting data amount. The settings follow Table 1.
Additionally, Fig. A3 shows the examples of generated images when forgetting the class ‘airplane’. It’s evident that opting for a 10% saliency sparsity can effectively forget the class ‘airplane,’ but it leads to a decrease in image generation quality. On the other hand, using a 90% saliency sparsity fails to forget the class. In Fig. A4, we present how the ‘optimal’ sparsity level can change vs. the number of forgetting data points, focusing on the case of MU for image classification. The optimal sparsity is determined through a greedy search with a sparsity interval of 10%. It’s important to note that the choice of a 50% sparsity ratio is not universal, and when dealing with a larger amount of forgetting data, a lower sparsity ratio is preferred. This is because a larger forgetting data set requires more weight updates, leading to a lower saliency sparsity.
@@ B.2 Additional classification results
Table A2: MU Performance on ResNet-18 for class-wise forgetting on CIFAR-10: Our method outperforms most approaches, although there is still a small gap with $\ell_{1}$-sparse in terms of UA.
@@ B.3 Additional generation results
Table A3: Class-wise forgetting on classifier-free guidance DDPM
Class UA ($\uparrow$) FID ($\downarrow$)
Retrain 98.20 20.50
ESD 100.00 47.57
SalUn 99.26 24.53
Extended from Fig. 4, We quantify the unlearning performance of Retrain, ESD, and SalUn in Table A3, using the two metrics introduced earlier, FID and UA. Comparing the performance of SalUn with ESD on DDPM with CIFAR-10, we observe a slight 0.84% UA drop in SalUn compared to ESD. However, SalUn significantly outperforms ESD in terms of FID. We notice that ESD exhibits instability when forgetting and learning low-quality images like CIFAR-10. Therefore, we believe that the 100% UA performance of ESD might also be due to the poor generation quality of images within the forgetting class. By contrast, SalUn yields the closest performance to Retrain. Fig. A5 and Fig. A6 shows the examples of generated images. The forgetting class is marked with a red border.
@@
In Fig. A8, we showcase the outcomes of class-wise unlearning on the ImageNette dataset utilizing the SalUn approach. The matrix configuration of the figure contrasts the “Unlearned class” with the “Prompt class”, distinctly delineating the desired versus the produced imagery. Images on the diagonal correspond to the target unlearning class, shedding light on the effectiveness of the SalUn method in this unlearning context. Conversely, off-diagonal images represent different classes, illustrating the model’s capability to differentiate and generalize across the broader dataset.
@@
Table A9 shows the prompts for Fig. 6.
@@ Appendix C Broader Impacts and Limitations
SalUn addresses some challenges in MU for Image Classification tasks, such as random data forgetting. For MU in Image Generation tasks, our method can effectively achieve both class-wise forgetting and concept-wise forgetting simultaneously. However, it’s worth noting that our approach may lead to a slight decrease in the quality of generated images. Furthermore, this paper primarily focuses on Computer Vision, and its effectiveness in Natural Language Processing still requires further exploration.
@@
Table A5: MU Performance across different forgetting data amounts on VGG-16, pre-trained on CIFAR-10 dataset, for random data forgetting. The content format follows Table 1.
@@
Table A6: MU Performance across different forgetting data amounts on Swin-T, pre-trained on CIFAR-10 dataset, for random data forgetting. The content format follows Table 1.

### 19 Feb 2024 v3

@@ 7 Conclusion
Recognizing the shortcomings of existing MU approaches, we introduce the innovative concept of weight saliency in MU, leading to the SalUn framework. This proposed saliency unlearning has showcased its effectiveness in addressing the limitations of current MU methods, and applying to both image classification and generation tasks. As a notable application, we show that SalUn stays effective in preventing stable diffusion from generating images with harmful content, even suffering inappropriate image prompts.
@@ 8 Acknowledgement
C. Fan, J. Liu, and S. Liu were supported by the Cisco Research Faculty Award and the National Science Foundation (NSF) Robust Intelligence (RI) Core Program Award IIS-2207052.
@@ Appendix A Soft-thresholding SalUn
where $i=1$ (or 22) corresponds to the objective function of SalUn for classification (6) (or generation (7)), and $\beta>0$ is a regularization parameter. Note that unlike (6) and (7), the objective function of SalUn is defined over the entire model weights $\bm{\theta}$} since there is no weight saliency map known in advance, i.e., letting $\mathbf{m}_{\mathrm{S}}=\mathbf{1}$ in (4). Problem (8) can be efficiently solved using the proximal gradient algorithm (Parikh et al., 2014), wherein the $\ell_{1}$ penalty term is handled efficiently through a closed-form proximal operation known as soft-thresholding. In contrast to the hard-thresholding implementation of SalUn, the soft-thresholding approach requires tuning an additional hyperparameter $\beta$, with the trade-off of not resorting to the hard threshold $\gamma$ in (3). In practice, we find that employing a linearly decaying scheduler for $\beta$ (i.e., promoting weight saliency sparsity during the early unlearning stages) yields better results than other schemes, approaching the performance of SalUn based on hard thresholding.
@@ B.1 Additional training and unlearning settings
For the Retrain method, training is conducted over $182$ epochs using the SGD optimizer with a cosine-scheduled learning rate initialized at $0.1$. For both FT and RL, training spans 10 epochs with a fixed learning rate of $0.1$. GA’s training settings involve a 5-epoch learning rate search within the interval $10^{-6},10^{-4}]$. In the case of IU, we explore the parameter $\alpha$ associated with the woodfisher Hessian Inverse approximation within the range $[1,20]$. For $\ell_{1}$-sparse, a learning rate search for the parameter $\gamma$ is executed within the range $10^{-6},10^{-4}]$, while maintaining a constant learning rate of 0.1. For the BS method, the FGSM step size is defined as 0.1. Both BS and BE methods involve a 10-epoch learning rate search in the interval $10^{-6},10^{-4}]$. Lastly, for SalUn and SalUn-soft, we trained for 10 epochs, searching for learning rates in the range $10^{-4},10^{-2}]$ and sparsity ratios in the range $[0.2,0.6]$.
@@
Refer to caption Refer to caption Refer to caption Refer to caption (a) UA for classification (b) MIA for classification (c) UA for generation (d) FID for generation
Figure A1: Performance of SalUn with different saliency sparsity vs Retrain. (a) and (b) use UA or MIA as metrics for classification. The settings follow Table 1. (c) and (d) use UA or FID as metrics for generation. The settings follow Fig 4. Points above Retrain indicate over-forgetting, while points below Retrain represent under-forgetting.
@@
Figure A2: Image generations of SalUn with different saliency sparsity, follows the same setting in Fig. 4.
Ablation study on sparsity choice of weight saliency.
Recall from (3) that the choice of the weight saliency sparsity threshold could be a crucial hyperparameter in our approach. In our experiments, we have set the default sparsity threshold to 50%. Fig. A1 provides a more detailed examination of the sparsity ratio. Specifically, Fig. A1-(a) and (b) present the performance of random data forgetting in image classification using ResNet-18 on CIFAR-10 (with the experiment setup same as Table 1). Fig. A1-(c) and (d) present the performance of class-wise forgetting for image generation using DDPM on CIFAR-10 with the same setting as Fig 4. In both cases, the performance of Retrain is provided for comparison. In the context of MU for image classification, choosing a 50% sparsity threshold is a reasonable option. A higher saliency sparsity may result in under-forgetting, as evidenced by the significant gap compared to Retrain as well as the lower unlearning accuracy (Jia et al., 2023) (namely, making it easier to classify the forgetting data points) or the lower MIA (Jia et al., 2023) (namely, making it challenging to infer the forgetting identity of a training point). This is not surprising since the higher saliency sparsity indicates fewer mode weights to be modified during unlearning. Conversely, selecting a lower sparsity ratio may result in over-forgetting due to the excessive number of weights to be modified. In the context of MU for image generation, a higher sparsity ratio also leads to under-forgetting, while a much lower sparsity ratio causes a rise in FID, introduced by over-forgetting.
Figure A3: The relationship between optimal saliency sparsity and forgetting data amount on random data forgetting under ResNet-18, CIFAR-10.
Additionally, Fig. A2 shows the examples of generated images when forgetting the class ‘airplane’. It’s evident that opting for a 10% saliency sparsity can effectively forget the class ‘airplane,’ but it leads to a decrease in image generation quality. On the other hand, using a 90% saliency sparsity fails to forget the class. In Fig. A3, we present how the ‘optimal’ sparsity level can change vs. the number of forgetting data points, focusing on the case of MU for image classification. The optimal sparsity is determined through a grid search with a sparsity interval of 10%. It’s important to note that the choice of a 50% sparsity ratio is not universal, and when dealing with a larger amount of forgetting data, a lower sparsity ratio is preferred. This is because a larger forgetting data set requires more weight updates, leading to a lower saliency sparsity.
@@ B.2 Additional classification results
Table A2: MU Performance for class-wise forgetting on ResNet-18, pre-trained on CIFAR-10 dataset. The content format follows Table 1.
@@
Table A3: Iterative Unlearning performance across different iteration on CIFAR-10 for random data forgetting. Forget 10% data each iteration. The content format follows Table 1.
Iteration # Methods Iterative Random Data Forgetting (10%, 5 Iterations)
UA RA TA MIA Avg. Gap
1 Retrain 5.24 100.00 94.26 12.88 0
FT 11.38 (6.14) 91.46 (8.54) 86.97 (7.29) 17.69 (4.81) 6.70
SalUn 1.82 (3.42) 99.84 (0.16) 94.09 (0.17) 14.47 (1.59) 1.34
2 Retrain 5.31 100.00 94.10 13.30 0
FT 10.60 (5.29) 97.27 (2.73) 87.81 (6.29) 19.42 (6.12) 5.11
SalUn 1.33 (3.98) 99.83 (0.17) 93.62 (0.48) 13.18 (0.12) 1.19
3 Retrain 6.64 100.00 92.78 14.60 0
FT 10.56 (3.92) 96.89 (3.11) 85.66 (7.12) 20.38 (5.78) 4.98
SalUn 2.07 (4.57) 99.90 (0.10) 93.43 (0.65) 22.13 (7.53) 3.21
4 Retrain 7.01 100.00 92.52 18.37 0
FT 8.82 (1.81) 97.64 (2.36) 85.42 (7.10) 18.62 (0.25) 2.88
SalUn 6.62 (0.39) 97.53 (2.47) 90.34 (2.18) 15.31 (3.06) 2.03
5 Retrain 7.91 100.00 91.72 19.29 0
FT 9.00 (1.09) 96.87 (3.13) 84.29 (7.43) 18.78 (0.51) 3.04
SalUn 6.63 (1.28) 99.42 (0.58) 91.62 (0.10) 26.00 (6.71) 2.17
Performance of iterative unlearning.
As illustrated in Table A3, we conduct iterative unlearning experiments by incrementally forgetting 10% of the data over five iterations (50% of data in total), i.e., for each iteration the forgetting set is 10% of the whole dataset, given ResNet-18 on the CIFAR-10 dataset. We evaluate the unlearning performance of our method by comparing it with the gold standard, Retrain. Additionally, we assess its performance in comparison to FT. Notably, even as data points are progressively forgotten, SalUn demonstrates a consistently minimal performance gap with Retrain, as evidenced by the smallest value in the Avg. Gap column.
@@
Table A4: MU Performance on ResNet-18, pre-trained on Tiny ImageNet dataset, for 10% random data forgetting.
Methods Random Data Forgetting (10%)
UA RA TA MIA Avg. Gap
Retrain 36.40 99.98 63.67 63.77 0
$\ell_{1}$-sparse 15.19 (21.21) 98.61 (1.37) 61.78 (1.89) 26.39 (37.38) 15.46
SalUn 17.37 (19.03) 98.73 (1.25) 61.67 (2.00) 66.90 (3.13) 6.35
Performance of Tiny ImageNet dataset
We conducted additional experiments on the Tiny ImageNet dataset (Le & Yang, 2015) with a higher resolution ($64\times 64$) than CIFAR-10 and CIFAR-100 in Table A4. We focused on evaluating our method against the baseline $\ell_{1}$-sparse and Retrain. Compared to $\ell_{1}$-sparse, SalUn demonstrates smaller gaps in terms of UA and MIA, with comparable RA and TA gaps. This shows an improved unlearning efficacy over $\ell_{1}$-sparse while preserving the model’s generalization ability post unlearning.
@@ B.3 Additional generation results
Table A5: Class-wise forgetting on classifier-free guidance DDPM.
Methods UA ($\uparrow$) FID ($\downarrow$)
Retrain 100.00 11.69
ESD 100.00 17.37
SalUn 100.00 11.21
Extended from Fig. 4, We quantify the unlearning performance of Retrain, ESD, and SalUn in Table A5, using the two metrics introduced earlier, FID and UA. Comparing the performance of SalUn with ESD on DDPM with CIFAR-10, we observe a slight 0.84% UA drop in SalUn compared to ESD. However, SalUn significantly outperforms ESD in terms of FID. We notice that ESD exhibits instability when forgetting and learning low-quality images like CIFAR-10. Therefore, we believe that the 100% UA performance of ESD might also be due to the poor generation quality of images within the forgetting class. By contrast, SalUn yields the closest performance to Retrain. Fig. A4-A6 shows the examples of generated images. The forgetting class is marked with a red border.
@@
In Fig. A7-A9, we showcase the outcomes of class-wise unlearning on the ImageNette dataset utilizing the SalUn approach under different random seeds. The matrix configuration of the figure contrasts the “Unlearned class” with the “Prompt class”, distinctly delineating the desired versus the produced imagery. Images on the diagonal correspond to the target unlearning class, shedding light on the effectiveness of the SalUn method in this unlearning context. Conversely, off-diagonal images represent different classes, illustrating the model’s capability to differentiate and generalize across the broader dataset.
@@
Table A11 shows the harmful text prompts for Fig. 6 respectively.
@@ Appendix C Broader Impacts and Limitations
SalUn marks an important advancement in addressing the multifaceted challenges of data privacy, security, and adherence to regulatory mandates. SalUn enhances unlearning effectiveness in machine learning models, maintaining their utility even under strict unlearning requirements. Its role in precluding the generation of harmful content underscores its capacity to foster societal norms and ethical standards. This proactive approach reduces the risk of generating inappropriate content also guides AI development towards alignment with ethical standards and societal expectations.
However, it’s crucial to acknowledge the limitations of our method. Although SalUn has proven effective in vision tasks, its scalability and adaptability to other domains like language and graphs require further investigation. The impact of machine unlearning on fairness, privacy, and security also demands careful consideration. Ensuring transparent, accountable, and inclusive development of these technologies is essential.
@@
Table A7: MU Performance across different forgetting data amounts on SVHN dataset for random data forgetting. The content format follows Table 1.
Methods Random Data Forgetting (10%)
UA RA TA MIA Avg. Gap RTE
Retrain 5.04 100.00 95.56 16.27 0 44.40
FT 0.44 (4.60) 100.00 (0.00) 95.78 (0.22) 2.62 (13.65) 4.62 2.76
RL 4.51 (0.53) 99.93 (0.07) 94.98 (0.58) 28.02 (11.75) 3.23 2.89
GA 0.67 (4.37) 99.54 (0.46) 95.52 (0.04) 1.16 (15.11) 4.99 0.11
IU 5.13 (0.09) 99.20 (0.80) 93.25 (2.31) 10.64 (5.63) 2.21 3.19
BE 0.42 (4.62) 99.54 (0.46) 95.69 (0.13) 1.16 (15.11) 5.08 0.20
BS 0.42 (4.62) 99.54 (0.46) 95.70 (0.14) 1.11 (15.16) 5.09 0.35
SalUn 5.04 (0.00) 99.55 (0.45) 94.16 (1.40) 15.89 (0.38) 0.56 2.91
SalUn-soft 5.31 (0.27) 99.41 (0.59) 94.21 (1.35) 14.07 (2.20) 1.10 2.98
Methods Random Data Forgetting (20%)
UA RA TA MIA Avg. Gap RTE
Retrain 4.88 100.00 95.50 15.66 0 41.01
FT 0.44 (4.44) 99.97 (0.03) 95.66 (0.16) 14.44 (1.22) 1.46 2.57
RL 2.63 (2.25) 99.89 (0.11) 95.14 (0.36) 16.50 (0.84) 0.89 2.90
GA 0.90 (3.98) 99.48 (0.52) 94.85 (0.65) 1.60 (14.06) 4.80 0.21
IU 2.88 (2.00) 97.22 (2.78) 91.28 (4.22) 8.37 (7.29) 4.07 3.23
BE 0.50 (4.38) 99.54 (0.46) 95.67 (0.17) 1.26 (14.40) 4.85 0.40
BS 0.51 (4.37) 99.55 (0.45) 95.68 (0.18) 1.26 (14.40) 4.85 0.73
SalUn 2.21 (2.67) 99.87 (0.13) 95.09 (0.41) 15.93 (0.27) 0.87 2.93
SalUn-soft 4.20 (0.68) 99.30 (0.70) 94.33 (1.17) 13.31 (2.35) 1.22 3.01
Methods Random Data Forgetting (30%)
UA RA TA MIA Avg. Gap RTE
Retrain 4.93 100.00 95.28 15.32 0 37.82
FT 0.44 (4.49) 100.00 (0.00) 95.71 (0.43) 2.34 (12.98) 4.47 2.36
RL 1.82 (3.11) 99.86 (0.14) 95.40 (0.12) 14.00 (1.32) 1.17 2.91
GA 3.30 (1.63) 96.76 (3.24) 89.90 (5.38) 9.65 (5.67) 3.98 0.30
IU 3.30 (1.63) 96.76 (3.24) 89.90 (5.38) 9.65 (5.67) 3.98 3.23
BE 0.46 (4.47) 99.54 (0.46) 95.69 (0.41) 1.96 (13.36) 4.68 0.59
BS 0.46 (4.47) 99.54 (0.46) 95.67 (0.39) 1.30 (14.02) 4.83 1.09
SalUn 4.60 (0.33) 99.12 (0.88) 94.41 (0.87) 14.00 (1.32) 0.85 2.93
SalUn-soft 4.83 (0.10) 99.30 (0.70) 94.46 (0.82) 13.83 (1.49) 0.78 2.97
Methods Random Data Forgetting (40%)
UA RA TA MIA Avg. Gap RTE
Retrain 5.06 100.00 95.22 16.64 0 34.39
FT 0.47 (4.59) 100.00 (0.00) 95.77 (0.55) 2.29 (14.35) 4.87 2.15
RL 2.21 (2.85) 99.79 (0.21) 95.13 (0.09) 13.81 (2.83) 1.50 2.89
GA 0.49 (4.57) 99.55 (0.45) 95.66 (0.44) 1.24 (15.40) 5.21 0.41
IU 1.71 (3.35) 98.55 (1.45) 91.93 (3.29) 6.51 (10.13) 4.55 3.21
BE 2.08 (2.98) 98.07 (1.93) 93.45 (1.77) 26.21 (9.57) 4.06 0.61
BS 0.48 (4.58) 99.55 (0.45) 95.64 (0.42) 1.61 (15.03) 5.12 1.43
SalUn 5.36 (0.30) 98.14 (1.86) 94.17 (1.05) 15.81 (0.83) 1.01 2.88
SalUn-soft 5.04 (0.02) 98.45 (1.55) 94.38 (0.84) 15.43 (1.21) 0.91 2.98
Methods Random Data Forgetting (50%)
UA RA TA MIA Avg. Gap RTE
Retrain 5.27 100.00 95.13 17.48 0 31.21
FT 0.46 (4.81) 100.00 (0.00) 95.78 (0.65) 2.33 (15.15) 5.15 1.95
RL 2.50 (2.77) 99.72 (0.28) 95.14 (0.01) 15.50 (1.98) 1.26 2.91
GA 0.46 (4.81) 99.54 (0.46) 95.70 (0.57) 1.20 (16.28) 5.53 0.50
IU 19.16 (13.89) 81.52 (18.48) 75.12 (20.01) 23.47 (5.99) 14.59 3.20
BE 35.52 (30.25) 65.00 (35.00) 59.65 (35.48) 56.54 (39.06) 34.95 0.82
BS 0.55 (4.72) 99.51 (0.49) 95.21 (0.08) 9.52 (7.96) 3.31 1.74
SalUn 6.03 (0.76) 97.72 (2.28) 93.66 (1.47) 17.52 (0.04) 1.14 2.92
SalUn-soft 5.14 (0.13) 98.73 (1.27) 93.97 (1.16) 15.58 (1.90) 1.11 2.90
Table A8: MU Performance across different forgetting data amounts on CIFAR-100 dataset for random data forgetting. The content format follows Table 1.
Methods Random Data Forgetting (10%)
UA RA TA MIA Avg. Gap RTE
Retrain 26.47 99.97 74.13 51.00 0 41.36
FT 2.42 (24.05) 99.95 (0.02) 75.55 (1.42) 11.04 (39.96) 16.36 2.27
RL 55.03 (28.56) 99.81 (0.16) 70.03 (4.09) 98.97 (47.97) 20.20 2.12
GA 3.13 (23.34) 97.33 (2.64) 75.31 (1.18) 7.24 (43.76) 17.73 0.13
IU 3.18 (23.29) 97.15 (2.82) 73.49 (0.64) 9.62 (41.38) 17.03 3.81
BE 2.31 (24.16) 97.27 (2.70) 73.93 (0.20) 9.62 (41.38) 17.11 0.25
BS 2.27 (24.20) 97.41 (2.56) 75.26 (1.13) 5.82 (45.18) 18.27 0.43
SalUn 27.58 (1.11) 99.80 (0.17) 70.80 (3.33) 94.06 (43.06) 11.92 2.13
SalUn-soft 24.24 (2.23) 98.95 (1.02) 70.48 (3.65) 79.13 (28.13) 8.76 2.54
Methods Random Data Forgetting (20%)
UA RA TA MIA Avg. Gap RTE
Retrain 26.84 99.99 73.88 52.41 0 36.88
FT 2.70 (24.14) 99.95 (0.04) 75.51 (1.63) 11.63 (40.78) 16.65 2.05
RL 54.74 (27.90) 99.47 (0.52) 65.59 (8.29) 97.32 (44.91) 20.41 2.11
GA 6.79 (20.05) 94.11 (5.88) 71.39 (2.49) 13.22 (39.19) 16.90 0.26
IU 5.34 (21.50) 95.54 (4.45) 70.89 (2.99) 11.79 (40.62) 17.39 3.77
BE 2.51 (24.33) 97.38 (2.61) 75.07 (1.19) 6.70 (45.71) 18.46 0.49
BS 2.53 (24.31) 97.38 (2.61) 75.05 (1.17) 6.57 (45.84) 18.48 0.82
SalUn 25.97 (0.87) 99.64 (0.35) 68.14 (5.74) 92.43 (40.02) 11.75 2.12
SalUn-soft 24.56 (2.28) 98.68 (1.31) 67.93 (5.95) 79.40 (26.99) 9.13 2.53
Methods Random Data Forgetting (30%)
UA RA TA MIA Avg. Gap RTE
Retrain 28.52 99.98 70.91 52.24 0 32.92
FT 2.65 (25.87) 99.94 (0.04) 75.17 (4.26) 11.18 (41.06) 17.81 1.44
RL 51.46 (22.94) 99.32 (0.66) 62.77 (8.14) 96.34 (44.10) 18.96 2.14
GA 2.40 (26.12) 97.39 (2.59) 75.33 (4.42) 5.70 (46.54) 19.92 0.40
IU 5.96 (22.56) 94.59 (5.39) 69.74 (1.17) 12.63 (39.61) 17.18 3.76
BE 2.44 (26.08) 97.37 (2.61) 74.77 (3.86) 6.53 (45.71) 19.56 0.76
BS 2.49 (26.03) 97.33 (2.65) 74.65 (3.74) 6.40 (45.84) 19.56 1.24
SalUn 24.51 (4.01) 99.53 (0.45) 66.27 (4.64) 90.01 (37.77) 11.72 2.16
SalUn-soft 27.21 (1.31) 97.96 (2.02) 64.79 (6.12) 78.15 (25.91) 8.84 2.56
Methods Random Data Forgetting (40%)
UA RA TA MIA Avg. Gap RTE
Retrain 30.07 99.99 69.87 58.06 0 28.29
FT 2.66 (27.41) 99.95 (0.04) 75.35 (5.48) 11.05 (47.01) 19.99 1.51
RL 51.75 (21.68) 99.27 (0.72) 59.41 (10.46) 95.78 (37.72) 17.64 2.12
GA 2.46 (27.61) 97.39 (2.60) 75.40 (5.53) 5.91 (52.15) 21.97 0.51
IU 4.58 (25.49) 96.29 (3.70) 70.92 (1.05) 10.32 (47.74) 19.49 3.78
BE 2.54 (27.53) 97.35 (2.64) 74.56 (4.69) 7.44 (50.62) 21.37 1.00
BS 2.70 (27.37) 97.26 (2.73) 74.10 (4.23) 7.63 (50.43) 21.19 1.66
SalUn 30.34 (0.27) 99.39 (0.60) 63.21 (6.66) 91.04 (32.98) 10.13 2.14
SalUn-soft 23.91 (6.16) 98.54 (1.45) 64.47 (5.40) 77.58 (19.52) 8.13 2.55
Methods Random Data Forgetting (50%)
UA RA TA MIA Avg. Gap RTE
Retrain 32.69 99.99 67.22 61.15 0 25.01
FT 2.71 (29.98) 99.96 (0.03) 75.11 (7.89) 10.71 (50.44) 22.08 1.25
RL 50.52 (17.83) 99.47 (0.52) 56.75 (10.47) 95.91 (34.76) 15.90 2.13
GA 2.61 (30.08) 97.49 (2.50) 75.27 (8.05) 5.92 (55.23) 23.97 0.66
IU 12.64 (20.05) 87.96 (12.03) 62.76 (4.46) 17.54 (43.61) 20.04 3.80
BE 2.76 (29.93) 97.39 (2.60) 74.05 (6.83) 8.85 (52.30) 22.92 1.26
BS 2.99 (29.70) 97.24 (2.75) 73.38 (6.16) 8.76 (52.39) 22.75 2.08
SalUn 29.15 (3.54) 99.38 (0.61) 61.20 (6.02) 89.94 (28.79) 9.74 2.13
SalUn-soft 23.26 (9.43) 98.32 (1.67) 63.08 (4.14) 77.90 (16.75) 8.00 2.50
Table A9: MU Performance across different forgetting data amounts on VGG-16 for random data forgetting. The content format follows Table 1.
@@
Table A10: MU Performance across different forgetting data amounts on Swin-T for random data forgetting. The content format follows Table 1.